# False Discovery Rate and Localizing Power

### v1

@@ 4.1.1 Correction strategies
In the absence of signal, that is, under the complete null hypothesis, and with independent test statistics, the canonical method worked well when false positives were counted among observed positive test statistics only, or among observed negative test statistics only. In these cases, the empirical false discovery proportion was within the confidence intervals, being neither conservative nor anti-conservative. When false positives were counted among all test statistics (both sides, positive and negative), the observed error rate was twice as large as for each side separately, and thus, twice as large as the nominal test level. Similar behavior was observed for the split-tails method. Combined and two-tailed methods controlled the error rate at the nominal level, albeit became conservative when only one or the other direction was considered.
With only positive or only negative true signal, all four correction strategies produced error rates that were within the expected level for test statistics with the same sign as the signal, even if leaning heavily towards conservativeness. However, for test statistics with the opposite sign as the signal, the observed proportion of false discoveries was not controlled by a large margin, approaching 100% error rate. When both sides were considered to count false positives (that is, all tests considered), the error rate was controlled within the nominal level, leaning again towards conservativeness, albeit less heavily.
With both positive and negative true signal in a balanced scenario, the error rate was controlled regardless of how the error rate was counted (positive test statistics only, negative only, or both). That control, however, leaned towards conservativeness. With unbalanced amounts of signal, which more typically represents real experiments, the error rate was not controlled in the side opposite to the one with a preponderance of true signal, with the nominal error rate lying outside the confidence interval.
@@ 4.1.3 Dependencies among tests
The same patterns above were observed for dependent tests. Introduced dependencies did not substantially alter the results when comparing the correction strategies or the different fdr methods.
@@ 4.2 Real data
Figure 3 shows maps thresholded at positive and negative sides using the different fdr methods and correction strategies, as well as with uncorrected p-values; the thresholds in each case are also shown.
Figure 3: Thresholded statistical maps using different fdr methods and correction strategies for directional inference, as well as uncorrected, for the narps data. Canonical and split-tails strategies produce generally similar results; likewise, combined and two-tailed methods produce broadly similar maps. Inferences using combined or two-tailed maps should not make directional statements about the results; these are only possible in the canonical and split-tails methods. The figure also shows slight but consistent gains in power with the bky procedure when compared to the usual bh procedure. Only visualization software that allow asymmetrical thresholds can be used to display fdr-corrected results that enable directional inferences.
@@ 5.1 Two-tailed tests
In effect, using two-tailed tests as default has at times been recommended for neuroimaging studies: Chen et al. (2019) argued that “it is difficult to imagine a scenario where this researcher would test all voxels simultaneously for just one directional change”. However, the researcher must be aware that the amount of errors is not guaranteed within a given direction if fdr is applied across all tests. Merely applying fdr methods as integrated into neuroimaging packages such as afni, fsl, FreeSurfer, or spm — each referenced in more than 1000 peer-reviewed publications annually (Poldrack et al., 2019) — then proceeding to directional inferences, does not control the error rate, as illustrated by the synthetic and real data.
@@ 5.2 Control of the error rate
The analyses using synthetic data indicate that, while fdr-based adjustment (with either bh or bky) controls the error rate in the set of tests to which the procedure is applied, should the researcher decide to make inferences about direction of effects, extra care is necessary, otherwise such inferences can be affected by excess error rate, which can well amount to up to 100% of the tests about which inference is made if true signal is present in only one side. Simple strategies such as using a two-tailed test, or combining the two sets of results with the directional hypotheses into one single fdr correction, do not guarantee control over the error rate for either direction, even though these control the error rate globally, i.e., when all tests are considered. Such global test, however, provides no information on the direction of effects.
Alternative approaches can be considered. For example, running fdr separately for the two set of p-values with each of the directional hypotheses (the canonical strategy), or running fdr separately for the positive and negative test statistics (the split-tails strategy). The analyses demonstrate that these approaches are able to control the error rate in each direction. However, they come with a penalty: should the null hypothesis be true everywhere, the error rate, which should be at the level $q$ under the complete null, is doubled, approaching $2q$.
@@
It should be understood, however, that areas of positive signal that appeared as significant with the two-tailed and combined strategies in the narps dataset example may have emerged simply because there is a preponderance of signal in the negative side of the map. This is because procedures that control fdr are adaptive, that is, they start with a stringent control for the smallest p-value, and proceed in steps, each contingent on the findings of the previous one, progressively loosening stringency and continuing as long as each subsequent test continue to be declared significant (Benjamini and Hochberg, 2000; Nichols, 2007). With two-tailed and combined strategies, the large amount of tests with strong signal in one side of the map (negative in our example) allows more opportunities for tests in the opposite side (positive) to be declared significant. The canonical and split-tails strategies break the opportunities for eventual strong or widespread signals present in one side impact the other.
@@ 5.3 Localizing power
Even though the above issues are a direct consequence of fdr not providing strong control over the error rate, routine strategies for data analysis and visualization may obscure this property. For example, the use of symmetrical thresholds is common in imaging visualization software, whereby the user loads a statistical map, and applies a threshold $t$; values above $+t$ are shown in a hot color scheme (e.g., red to yellow), while values below $-t$ are shown in a cool color scheme (e.g., blue to cyan); for two-tailed tests, this seems appropriate and rigorous. Yet, the amount of errors in any direction is not controlled with fdr, only the global amount of errors, something not apparent to the user of such tools. Another situation is when drawing conclusions about the amount of errors within a region after controlling globally; it might seem that controlling the error rate for the whole brain is the most rigorous, such that inferences within sub-regions would be conservative, given that a larger number of tests is taken into account. Yet, the amount of errors within any particular region is not controlled, only the global amount of errors. Any attempt to isolate a particular direction or a particular subset of tests will fail to guarantee the same proportion $q$ of false discoveries within that subset. This example also makes it clear that fdr does not provide localizing power; the proportion of errors is guaranteed only globally.
As mentioned in Section 2.8, such localization and direction errors can be avoided by refraining from making localized or directional statements encompassing subsets of the tests, making only such statements when they refer to the whole set of tests that were subjected to the fdr procedure. Direction errors can also be avoided by using strategies such as those that we named canonical or split-tails. These approaches, however, require that the user is able to display the statistical maps in software that enables non-symmetric thresholds to be used. That is, the ideal imaging tool should accommodate separate thresholds for positive and for the negative side of the map, such that these can be set individually after running fdr with one of these strategies. Currently, very few software packages allow the user to easily set asymmetric thresholds for positive and negative sides of the map. Having this feature available across different imaging software would more easily allow for directional investigations with fdr and visualization of results, even using two-tailed tests.
@@ 5.4 More power with bky
An important yet little appreciated fact is that the bh procedure — the most commonly adopted — does not control the amount of false discoveries at the exact level $q$ desired by the researcher, but at a lower level $qV_{0}/V$ that is determined by the (unknown) proportion of tests in which the null hypothesis is true among the total number of tests (Section 2.1). Hence, the procedure is, in fact, conservative in relation to the level desired by the researcher. Despite this, it remains more powerful than procedures that control fwer such as Šidák or Bonferroni; typically, unless $V_{0}=V$ (complete null), $\textsc{fwer}=1$ for any fdr-controlling procedure. Whereas bh is adaptive in the sense that the threshold that determines whether a p-value is significant varies according to the sorted p-value position $i$, with fixed slope given by $q/V$, fwer-controlling procedures as Bonferroni use the same constant, horizontal threshold for all p-values. The bky procedure, allows one to eke out more power by being even more adaptive than bh in that the slope itself, rather than fixed, varies according to the sorted p-value position $i$ as $q/(V+1-i(1-q))$. Thus, as far as power considerations are concerned, the bky procedure is preferable over the bh procedure.
This extra power comes with a computational price: bh runs in $\mathcal{O}(V)$ time, whereas bky runs in $\mathcal{O}(V^{2})$ time. This means that, while the time it takes to run fdr grows linearly with the number of tests to be corrected, for bky it grows quadratically. Another point worth mentioning is that bky-adjusted p-values can easily be above 1 (Figure 2. This is not a unique feature of bky: other multiple testing procedures can also lead to adjusted p-values being higher than 1, the simplest being Bonferroni, that leads to adjusted p-values that are equal to the uncorrected p-values multiplied by the number of tests $V$. This fact does not have substantial implications for hypotheses testing, and can be solved via Winsorization at 1.
@@ 6 Conclusion
While fdr procedures provide more power than fwer methods if many tests are expected to be significant and the user is willing to accept some false positives, caution is warranted when making directional inferences after using two-tailed tests. Strategies like running fdr separately on positive and negative test statistics, or on one-tailed p-values partitioned by hypothesis direction, avoid direction errors and control the false discovery proportion in each tail. However, these strategies lose fdr benefits under the complete null, such as controlling the fwer in these cases. Researchers should refrain from making directional inferences about fdr-corrected maps after two-tailed tests particularly when signals are not symmetrically distributed, since weak control means the proportion of false discoveries is only guaranteed globally; for the same reason, localization of effects within subsets is not possible. The bky procedure allows for additional power, albeit it does not eliminate issues with asymmetric signals. Allowing asymmetric thresholds in neuroimaging software would enable appropriate visualization after partitioned fdr.
@@ Acknowledgements
All authors receive support from the National Institutes of Health. A.M.W.: U54-HG013247; P.T.: ZIC-MH002888; T.E.N.: R01-EB026859, U24-DA041123, and U19-AG073585; C.R.: RF1-MH133701 and P50-DC014664.

### v2

@@ 4.1.1 Correction strategies
In the absence of signal, that is, under the complete null hypothesis, and with independent test statistics (scenario i), the canonical method worked well when false positives were counted among observed positive test statistics only, or among observed negative test statistics only. In these cases, the empirical false discovery proportion was within the confidence intervals, being neither conservative nor anti-conservative. When false positives were counted among all test statistics (both sides, positive and negative), the observed error rate was twice as large as for each side separately, and thus, twice as large as the nominal test level. Similar behavior was observed for the split-tails method. Combined and two-tailed methods controlled the error rate at the nominal level, albeit became conservative when only one or the other direction was considered.
With only positive or only negative true signal (scenarios ii and iii), all four correction strategies produced error rates that were within the expected level for test statistics with the same sign as the signal, even if leaning heavily towards conservativeness. However, for test statistics with the opposite sign as the signal, the observed proportion of false discoveries was not controlled by a large margin, approaching 100% error rate. When both sides were considered to count false positives (that is, all tests considered), the error rate was controlled within the nominal level, leaning again towards conservativeness, albeit less heavily.
With both positive and negative true signal in a balanced scenario (scenario iv), the error rate was controlled regardless of how the error rate was counted (positive test statistics only, negative only, or both). That control, however, leaned towards conservativeness. With unbalanced amounts of signal (scenario v), which more typically represents real experiments, the error rate was not controlled in the side opposite to the one with a preponderance of true signal, with the nominal error rate lying outside the confidence interval, and conservative on the opposite side.
@@ 4.1.3 Dependencies among tests
The same patterns above were observed for dependent tests (scenarios vix). Introduced dependencies did not substantially alter the results when comparing the correction strategies or the different fdr methods.
@@ 4.2 Real data
Figure 3 shows maps thresholded at positive and negative sides using the different fdr methods and correction strategies, as well as with uncorrected p-values; the thresholds in each case ($t$ statistics) are also shown.
Figure 3: Thresholded statistical maps using different fdr methods and correction strategies for directional inference, as well as uncorrected, for the narps data. Significant results in the positive side are those greater than the threshold $t_{\text{pos}}$; in the negative side, those smaller (more negative) than $t_{\text{neg}}$; if nothing is declared significant, the threshold is set, for display purposes, at a value more extreme than the test statistic at that side, here generically represented as $\infty$. The canonical and split-tails strategies produce generally similar results; likewise, combined and two-tailed methods produce broadly similar maps. Inferences using combined or two-tailed maps should not make directional statements about the results; these are only possible in the canonical and split-tails methods. The figure also shows slight but consistent gains in power with the bky procedure when compared to the usual bh procedure. Only visualization software that allow asymmetrical thresholds can be used to display fdr-corrected results that enable directional inferences.
@@ 5.1 Two-tailed tests
In effect, using two-tailed tests as default has at times been recommended for neuroimaging studies: Chen et al. (2019) argued that “it is difficult to imagine a scenario where this researcher would test all voxels simultaneously for just one directional change”. However, the researcher must be aware that the amount of errors is not guaranteed within a given direction if fdr is applied across all tests. Merely applying fdr methods as integrated into neuroimaging packages such as afni, fsl, FreeSurfer, or spm — each referenced in thousands of peer-reviewed publications annually (Poldrack et al., 2019) — and then proceeding to directional inferences, does not control the error rate within each direction, as illustrated by the synthetic and real data.
@@ 5.2 Control of the error rate
The analyses using synthetic data indicate that, while fdr-based adjustment (with either bh or bky) controls the error rate in the set of tests to which the procedure is applied, should the researcher decide to make inferences about direction of effects, extra care is necessary, otherwise such inferences can be affected by excess error rate, which can well amount to up to 100% of the tests about which inference is made if true signal is present in only one side. Simple strategies such as using a two-tailed test followed by fdr, or combining the two sets of results with the directional hypotheses into one single fdr correction, do not guarantee control over the error rate for either direction, even though these control the error rate globally, i.e., when all tests are considered. Such global test, however, provides no information on the direction of effects.
Alternative approaches were considered here. For example, running fdr separately for the two set of p-values with each of the directional hypotheses (the canonical strategy), or running fdr separately for the positive and negative test statistics (the split-tails strategy). The analyses demonstrate that these approaches are able to control the error rate in each direction. However, they come with an important penalty: should the null hypothesis be true everywhere, the error rate, which should be at the level $q$ under the complete null across both sides, is doubled, approaching $2q$ when both sides are considered. Should the researcher be interested in retaining the error controlled even in the complete null, then the test level should be halved. It should be noted that trying to control fdr for all subsets simultaneously would transform it into an fwer-controlling procedure (Finner and Roters, 2001).
@@
It should be understood, however, that areas of positive signal that appeared as significant with the two-tailed and combined strategies in the narps dataset example may have emerged simply because there is a preponderance of signal in the negative side of the map; evidence for that comes from the results using synthetic data. This is because procedures that control fdr are adaptive, that is, they start with a stringent control for the smallest p-value, and proceed in steps, each contingent on the findings of the previous one, progressively loosening stringency and continuing as long as each subsequent test continue to be declared significant (Benjamini and Hochberg, 2000; Nichols, 2007). With two-tailed and combined strategies, the large amount of tests with strong signal in one side of the map (negative in our example) allows more opportunities for tests in the opposite side (positive) to be declared significant. The canonical and split-tails strategies eliminate the opportunity for eventual strong or widespread signals present in one side impact the other, as fdr is applied separately to each side.
@@ 5.3 Localizing power
Even though the above issues are a direct consequence of fdr not possessing the subsetting property, i.e., not providing localizing power, routine strategies for data analysis and visualization may obscure this property. For example, the use of symmetrical thresholds is common in imaging visualization software: the user loads a statistical map and applies a threshold $t$; values above +t+t are shown in a hot color scheme (e.g., red to yellow), while values below t-t are shown in a cool color scheme (e.g., blue to cyan); for two-tailed tests, this seems appropriate and rigorous. Yet, the amount of errors in any direction is not controlled with fdr, only the global amount of errors, something that is not always apparent to the user of such tools, or they may (incorrectly) assume that the fdr is controlled in the positive and negative sides of the map. Another situation is when drawing conclusions about the amount of errors within a region after controlling globally; it might seem that controlling the error rate for the whole brain is the most rigorous, such that inferences within sub-regions would be conservative, given that a larger number of tests is taken into account. Yet, the amount of errors within any particular region is not controlled, only the global amount of errors. Any attempt to isolate a particular direction or a particular subset of tests will fail to guarantee the desired proportion of false discoveries within that subset.
@@ 5.4 More power with bky
An important yet little appreciated fact is that the bh procedure — the most commonly adopted — does not control the amount of false discoveries at the exact level $q$ desired by the researcher, but at a lower level $qV_{0}/V$ that is determined by the (unknown) proportion of tests in which the null hypothesis is true among the total number of tests (Section 2.3). Hence, the procedure is, in fact, conservative in relation to the level desired by the researcher. Despite this, it remains more powerful than procedures that control fwer such as Šidák or Bonferroni that use, for all tests, a fixed significance threshold that determines whether a p-value is significant; unless $V_{0}=V$ (complete null), $\textsc{fwer}=1$ for any fdr-controlling procedure. Whereas bh is adaptive in the sense that the threshold varies according to the sorted p-value at position $i$, with a fixed slope given by $q/V$, the bky procedure is even more adaptive than bh in that the slope itself varies according to the sorted p-value at position $i$ as q/(V+1i(1q))q/(V+1-i(1-q)) (Figure 1). The quantity q/(V+1i(1q))q/(V+1-i(1-q)) is only smaller than $q/V$ if i<1/(1q)i<1/(1-q). Thus, for typical $q$, usually well below 0.5, bky is more powerful than bh as long as smallest ($i=1$) test is rejected. As far as power considerations are concerned, the bky procedure is therefore preferable over the bh procedure.
This extra power comes with a computational price: bh runs in $\mathcal{O}(V)$ time, whereas bky runs in $\mathcal{O}(V^{2})$ time. This means that, while the time it takes to run fdr grows linearly with the number of tests to be corrected, for bky it grows quadratically. Another point worth mentioning is that bky-adjusted p-values can easily be above 1 (Figure 2). This is not a unique feature of bky: other multiple testing procedures can also lead to adjusted p-values being higher than 1, the simplest being Bonferroni, that leads to adjusted p-values that are equal to the uncorrected p-values multiplied by the number of tests $V$. This fact does not have substantial implications for hypotheses testing, and can be solved via truncation at 1.
The bky procedure was developed as an improvement over bh, and thus it would seem that the same considerations regarding dependencies between tests would hold. However, results have only been proven for the independent case (Benjamini et al., 2006), even if it tends to perform well in simulations under certain forms of dependence (e.g., Stevens et al., 2017). Scaling the thresholds of bky by $1/c(V)$ to accommodate arbitrary dependencies, as in by, while compelling, has not been proven to control the fdr either (nor that it does not).
@@ 5.5 Dependencies among tests
5.6 Visualization of results
As mentioned in Section 2.6 and 5.3, localization and direction errors can be avoided by refraining from making localized or directional statements encompassing subsets of the tests, making only such statements when they refer to the whole set of tests that were subjected to the fdr procedure. Direction errors can also be avoided by using strategies such as those that we named canonical or split-tails. These approaches, however, require that the user is able to display the statistical maps in software that enables non-symmetric thresholds to be used. That is, the ideal imaging tool would accommodate separate thresholds for positive and for the negative sides of the map, such that these can be set individually after running fdr with one of these strategies. This can work also with approaches that do not use abrupt thresholds, but instead a transparency gradient, allowing subthreshold evidence be viewed informatively while highlighting suprathreshold results (Allen et al., 2012; Taylor et al., 2025).
Currently, very few software packages allow the user to easily set asymmetric thresholds for positive and negative sides of the map. Having this feature available across different imaging software would more easily allow for directional investigations with fdr and visualization of results, even using two-tailed tests. For example, a map of uncorrected p-values can be thresholded at the fdr-controlling level, or a map of adjusted p-values can be thresholded at the desired level. Alternatively, if the distribution of the test statistic is known, the threshold p-value can be converted to a threshold statistic (using the cumulative distribution function, parametric or non-parametric), which can be applied to partition statistical map into significant and non-significant regions. A thresholded map (of statistics or of p-values) can be binarized and used as a mask to display or select regions for further analyses; the same workaround can be used for multiple regions, each of them subjected to correction separately, possibly with the test level further scaled down by the number of such regions should the researcher wish to control the error rate under the complete null.
5.7 Recommendations
When applying the fdr in neuroimaging studies with two-tailed tests, researchers should adapt their approach to their goals. fdr does not ensure error control within subsets, so localized inferences require applying fdr separately, possibly with the test level adjusted to accommodate the case of no effects anywhere. For directional inferences, such as detecting positive or negative effects, the canonical or split-tails strategies are recommended to control the fdr within each direction, though the total fdr may double in the complete absence of effects; the test level is proportionally reduced if that is a concern. To apply this approach, software have to allow thresholding to be asymmetric. For non-directional inferences, the two-tailed or combined strategies suffice. The bky procedure offers greater power than the bh method.
@@ 6 Conclusion
While fdr procedures provide more power than fwer methods if many tests are expected to be significant and the user is willing to accept some false positives, caution is warranted when making directional inferences after using two-tailed tests. Strategies like running fdr separately on positive and negative test statistics, or on one-tailed p-values partitioned by hypothesis direction, avoid direction errors and control the false discovery proportion in each tail. However, these strategies lose fdr benefits under the complete null, such as controlling the fwer in these cases. Researchers should refrain from making directional inferences about fdr-corrected maps after two-tailed tests particularly when signals are not symmetrically distributed, since weak control means the proportion of false discoveries is only guaranteed globally; for the same reason, localization of effects within subsets is not possible using these methods. The bky procedure allows for additional power, albeit it does not eliminate issues with asymmetric signals. Allowing asymmetric thresholds in neuroimaging software would enable appropriate visualization after partitioned fdr.
@@ Acknowledgements
Support was provided by the National Institutes of Health. A.M.W.: U54-HG013247; P.A.T.: ZIC-MH002888; T.E.N.: R01-EB026859, U24-DA041123, and U19-AG073585; C.R.: RF1-MH133701 and P50-DC014664. The authors are grateful to the anonymous reviewers who provided helpful feedback.